Question 10

A Firm's marginal cost function is Q2 + 2Q + 4


Find the total cost function.


a. Q3/3 + Q2 + 4 + C

b. Q3/3 + Q2 + C

c. Q3/3 + Q2 + 4Q + C

d. Q3/3 + Q2 + 100


1
Expert's answer
2021-06-24T06:42:36-0400
"MC=Q^2+2Q+4"

"TC=\int MCdQ=\int (Q^2+2Q+4)dQ"

"=\dfrac{Q^3}{3}+Q^2+4Q+C"

c. "\dfrac{Q^3}{3}+Q^2+4Q+C"